# Constrained Optimization From a Control Perspective via Feedback Linearization

### v2

@@ 3.2 Relationship with SQP
$\dot{x}\!=\!-T(x)\left(\nabla f(x)\!-\!\!J_{h}(x)^{\top}\left({J_{h}(x)T(x)J_{h}(x)^{\top}}\right)^{\!-\!1}\cdot(J_{h}(x)T(x)\nabla f(x)\!-\!Kh(x))\right).$
@@
$x_{t+1}=x_{t}-\eta T(x_{t})\left(\nabla f(x_{t})-J_{h}(x_{t})^{\!\top}\!\cdot\!\left(\!J_{h}(x_{t})T(x_{t})J_{h}(x_{t})^{\!\top}\!\right)^{\!-\!1}\!\!(J_{h}(x_{t})T(x_{t})\nabla\!f(x_{t})\!-\!Kh(x_{t}))\!\right)\!.$ (7)
@@
$\!x_{\!t\!+\!1}$ $\!=\!\operatorname*{arg\,min}_{x}\nabla\!f(x_{t})^{\!\top}\!(x\!-\!x_{t})\!+\!\frac{1}{2\eta}\!(x\!-\!x_{t})^{\!\top}\!T(x_{t})^{-1}(x\!-\!x_{t})$
$\qquad s.t.\quad h(x_{t})+J_{h}(x_{t})(x-x_{t})=0$ (8)
@@
When $K=\frac{1}{\eta}I$, the discretization of feedback linearization (7) is equivalent to the SQP algorithm (3.2).
The proof of Theorem 2 leverages the fact that (3.2) satisfies the relaxed Slater condition, and thus the KKT conditions are necessary and sufficient for global optimality. Then Theorem 2 can be obtained by studying the KKT conditions of 3.2. The detailed proof is deferred to Appendix C
@@
Theorem 2 provides insights into the selection of $T(x)$ for the FL approach. Different choices of $T(x)$ will correspond to different types of SQP algorithms. Here we mainly discuss two specific types of $T(x)$. Firstly, when $T(x)$ is chosen as the inverse of the Hessian matrix, i.e., $T(x)=\left(\nabla^{2}f(x)\right)^{-1}$, then (3.2) corresponds to the Newton-type algorithm where the quadratic term in the objective function is given by $(x-x_{t})^{\top}\nabla^{2}f(x)(x-x_{t})$, which is widely considered in literature (cf. Nocedal and Wright (2006); Bonnans et al. (2006)). For this specific type of $T(x)$, we name its corresponding FL dynamics (5) as the FL-Newton method. However, in the setting where the Hessian information is not available, another choice of $T(x)$ is simply setting it as the identity matrix $T(x)=I$, which is considered in recent works such as Oztoprak et al. (2021). In this case, the objective function resembles a proximal operator (cf. Boyd (2004); Parikh et al. (2014)), hence we name this as FL-proximal method. Due to space limit, we defer a more comprehensive overview of SQP in to Appendix A.
@@ 4 Extension to inequality constraints
$\vskip-5.0pt\begin{split}-\nabla f(x)-J_{h}(x)^{\top}\lambda=0,\quad h(x)\leq 0\\ \lambda\geq 0,\quad\lambda^{\top}h(x)=0\end{split}\vskip-5.0pt$ (10)
Thus, we can still view the problem as a control problem whose corresponding dynamics can be written as:
$\begin{split}\dot{x}&=-T(x)\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)\\ y&=h(x),\quad\lambda\geq 0.\end{split}$ (11)
However, the problem becomes more complicated because we require the non-negativity constraints $\lambda\geq 0$ and complementary slackness $\lambda^{\top}h(x)=0$. It is at first glance unclear how to guarantee theses conditions through the control process. However, inspired by the relationship with SQP algorithms, we carefully design a more intricate FL controller as follows:
FL for Inequality-Constrained Optimization $\begin{split}\dot{x}&\!=\!-T(x)\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)\\ \lambda&=\operatorname*{arg\,min}_{\lambda\geq 0}\!\left(\frac{1}{2}\lambda^{\!\top}{J_{h}(x)T(x)J_{h}(x)^{\top}}\lambda+\lambda^{\top}\left(J_{h}(x)T(x)\nabla f(x)\!-\!Kh(x)\right)\right)\end{split}$ (12)
Here we assume that the optimization problem
$\lambda$ $=\operatorname*{arg\,min}_{\lambda\geq 0}\!\left(\frac{1}{2}\lambda^{\!\top}{J_{h}(x)T(x)J_{h}(x)^{\top}}\lambda+\lambda^{\top}\left(J_{h}(x)T(x)\nabla f(x)\!-\!Kh(x)\right)\right)$
admits a unique solution. We would also like to point out that $\lambda$ in (12) takes the form of the solution of an optimization problem, resulting in a non-smooth trajectory. A similar formulation of non-smooth ordinary differential equations (ODEs) has been explored in the context of differential variational inequalities (cf. Dupuis and Nagurney (1993); Pang and Stewart (2008); Camlibel et al. (2007)).
@@
Similar to the proof of Theorem 2, the proof of Theorem 3 also leverages strong duality and KKT conditions. THe detailed proof is deferred to Appendix C.
@@
We define the index set $\mathcal{I}(x):=\{i:h_{i}(x)>0\}$. We also use $\mathcal{I}(x)^{c}$ to denote the complimentary set of $\mathcal{I}(x)$. Our results rely on the following assumptions:
@@
We make the following assumptions on the function $f$ and $h$
  1. 2.1
    Given the initial state $x(0)$ at $t=0$, the optimization problem in (12)
    $\vskip-5.0pt\lambda$ $=\textstyle\operatorname*{arg\,min}_{\lambda\geq 0}\!\Big{(}\frac{1}{2}\lambda^{\!\top}{J_{h}(x)T(x)J_{h}(x)^{\top}}\lambda+\lambda^{\top}\left(J_{h}(x)T(x)\nabla f(x)\!-\!Kh(x)\right)\Big{)}$
    admits bounded a solution $\|\lambda\|_{\infty}\leq L$ for all $x\in\mathcal{E}$, where $\mathcal{E}$ is defined by $\mathcal{E}:=\{x|0<h_{i}(x)\leq h_{i}(x(0)),~\forall i\in\mathcal{I}(x(0))\}$.
  2. 2.2
    There exists a constant $M$ such that $\|\nabla f(x)\|<M$, $\|J_{h}(x)\|<M$ for all $x$.
  3. 2.3
    The function $f(x)$ is lower-bounded, i.e. $f(x)\geq f_{\min}$ for all $x$.
Although Assumption 2.1 is quite complicated and relatively hard to verify, there are some simplified versions that serve as a sufficient condition of Assumption 2.1. For example, if we start with a feasible $x(0)$, then $\mathcal{E}=\emptyset$ and hence Assumption 2.1 is automatically satisfied. Additionally, note that Assumption 1.1 is another sufficient condition of Assumption 2.1 (see Lemma 3 in Appendix F)
We define the KKT-gap of the state variable $x$ and nonnegative control variable $\lambda\geq 0$ as follows:
$\textup{{KKT-gap}}(x,\lambda):=\max\left\{\|\nabla f(x)+J_{h}(x)^{\top}\lambda\|,\left|\lambda^{\top}h(x)\right|,\max_{i}[h_{i}(x)]_{+}\right\},$
where $h_{i}(x)]_{+}=\max\{h_{i}(x),0\}$.
@@
Under Assumption 2, for a diagonal matrix $K=\textup{diag}\{k_{i}\}_{i=1},$ where $k_{i}>0$, the learning dynamics (12) satisfies the following properties
@@
    $\frac{dh_{i}(x(t))}{dt}\leq-k_{i}h_{i}(x(t))$, for $i=1,2,\dots,m$, and hence the dynamic will asymptotically converge to the feasible set.
@@
    $\hskip-16.0pt\int_{t\!=\!0}^{T}\!\!\left(\!\!\|\nabla f(x(t))\!+\!\!J_{h}(x(t))\lambda(t)\|_{T(x(t))}^{2}\!\!-\!\!\!\!\!\sum_{i\in\!\mathcal{I}(x)^{c}}\!\!\!k_{i}\lambda_{i}(t)h_{i}(x(t))\!\!\right)\!\!dt$
    $\leq\!\ell(x(0))\!-\!\ell(x(T))$
@@
    $\inf_{0\leq t\leq T}\textup{{KKT-gap}}(x(t),\lambda(t))$
    $~~~\left.\frac{1}{\min_{i}k_{i}}\frac{2}{T}\left(\!f(x(0))\!-\!f_{\min}\!+\!(L\!+\!1)\!\!\!\!\sum_{i\in\mathcal{I}(x(0))}\!\!\!\!h_{i}(x(0))\!\right)\!\right\}$
    $\sim O\left(\frac{1}{\sqrt{T}}\right)$
    Further we have that KKT-gap asymptotically converges to zero, i.e.
    $\lim_{t\to+\infty}\textup{{KKT-gap}}(x(t),\lambda(t))=0$
Statement 4 in Theorem 4 implies that the algorithm can find an $\epsilon$-firs-order KKT-point within time $\frac{1}{\epsilon^{2}}$. Similar to Theorem 1, the key step of the proof is to construct the Lyapunov function in Statement 2 (detailed proof deferred to Appendix D).
@@ 5 Momentum Acceleration for Constrained Optimization
$\vskip-5.0pt\begin{split}w_{t}&=x_{t}+\beta(x_{t}-x_{t-1})\\ x_{t+1}&=w_{t}-\eta\nabla f(w_{t})\end{split}\vskip-5.0pt$ (15)
Its corresponding continuous-time analogue can be written as a second-order ODE Polyak (1964); Su et al. (2016):
[…]
@@
$w_{t}=x_{t}+\beta(x_{t}-x_{t-1})$
$\lambda_{t}\!=\!-\!\left(J_{h}(w_{t})J_{h}(w_{t})^{\!\top}\right)^{\!-\!1}\!\!(J_{h}(w_{t})\nabla f(w_{t})\!-\!\frac{1}{\eta}h(w_{t}))$
$x_{t+1}=w_{t}+\eta\nabla f(w_{t})+J_{h}(w_{t})^{\top}\lambda_{t}$ (17)
@@
FL-momentum for Equality-Constrained Optimization $\begin{split}\dot{x}&=z\\ \dot{z}&=-\alpha z-\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)\\ \lambda&=-(J_{h}(x)J_{h}(x)^{\top}\!)^{\!-1}(J_{h}(x)\nabla f(x)\!-\!Kh(x))\end{split}$ (18)
Note that the only difference in (5) is that we add a momentum step $w_{t}=x_{t}+\beta(x_{t}-x_{t-1})$. Similarly we can propose the FL-momentum scheme for inequality constraint case as follows:
FL-momentum for Inquality-Constrained Optimization $\dot{x}$ $=z$ $\dot{z}$ $=-\alpha z-\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)$ (19) $\lambda$ $\!=\!\operatorname*{arg\,min}_{\lambda\geq 0}\!\frac{1}{2}\lambda^{\!\top}\!J_{h}(x)J_{h}(x)^{\!\top}\!\lambda\!+\!\lambda^{\!\top}\!\!\left(J_{h}(x)\nabla\!f(x)\!-\!Kh(x)\right)$
The numerical simulation in Section 6 (Figure 2 and 3) demonstrates the comparison between the standard and momentum accelerated methods, which suggests that momentum methods indeed accelerate the convergence rate.
We would also like to note that as far as we know, the acceleration of SQP methods are generally achieved via Newton or quasi-Newton methods, there’s little work on exploring acceleration via momentum approaches, which makes our proposed momentum algorithm a novel contribution.
5.1 Analysis
@@
Apart from Assumption 1, we also make the following assumptions on $f$ and $h$.
@@
Both $f(x)$, $h(x)$ are three-time differentiable and the derivatives are bounded, thus, we know that there exists some constant $L_{f},L_{1},L_{2}$ such that
$\qquad\|\nabla^{2}f(x)\|\leq L_{f},~~\left\|\frac{\partial\bar{\lambda}(x)}{\partial x}\right\|\leq L_{2},$
$\left\|\frac{\partial\left(J_{h}(x)^{\top}\lambda(x)+\left(\frac{\partial\bar{\lambda}(x)}{\partial x}^{\top}h(x)\right)\right)}{\partial x}\right\|\leq L_{2}$
@@
We also assume that that there exists a constant $\bar{H}$ such that
$\|\bar{H}(x)\|\leq\bar{H},~~\forall x,$
where $\bar{H}(x):=[h(x)^{\top}\nabla^{2}h_{i}(x)]_{i=1}^{n}$.
@@
Assume that Assumption 1, 3 and 4 hold. Let two positive constant $a_{1},a_{2}$ be such that
$\textstyle a_{2}\geq\left(4\frac{\lambda_{\max}(K)}{{\lambda_{\min}}(K)}L_{2}D+\frac{L_{1}^{2}}{{\lambda_{\min}}(K)}\right)\times a_{1}\geq 0.$
We define the following Lyapunov function:
@@
then for
we have that
@@
    $\hskip-20.0pt\int_{t\!=\!0}^{T}\!\!\!\frac{a_{2}{\lambda_{\min}}(K)}{8}\|h(x(t))\|^{2}\!\!+\!\!\frac{a_{1}}{4}\|\nabla\!f(x(t))\!\!+\!\!J_{h}(x(t))^{\top}\overline{\lambda}(x(t))\|^{2}dt$
    $\leq\ell(x(0),z(0))-\min_{x,z}\ell(x,z)$
@@
    $\quad\inf_{0\leq t\leq T}\textup{{KKT-gap}}(x(t),\overline{\lambda}(x(t)))$
    $\leq\sqrt{\frac{\ell(x(0),z(0))-\ell_{\min}}{\min\left\{\frac{a_{2}{\lambda_{\min}}(K)}{8},\frac{a_{1}}{4}\right\}T}}\sim O(\frac{1}{\sqrt{T}})$
    and
    $\textstyle\lim_{t\to+\infty}\textup{{KKT-gap}}(x(t),\overline{\lambda}(x(t)))=0,$
    $\textstyle\lim_{t\to+\infty}\textup{{KKT-gap}}(x(t),\lambda(t))=0.$
The detailed proof is provided in Appendix E. One limitation of Theorem 5 is that it establishes convergence but not acceleration over FL-proximal. However, when the constraint function $h(x)$ is affine, the algorithm is equivalent to the momentum-accelerated projected gradient method (see Appendix E.1), offering insight into its potential for accelerating optimization.
6 Numerical Simulation
6.1 Heterogeneous Logistic Regression
In this section, we consider a logistic regression problem involving heterogeneous clients Shen et al. (2022); Hounie et al. (2024). Many scenarios, such as federated learning and fair machine learning, require training a common model in a distributed manner by utilizing data samples from diverse clients or distributions. In practice, heterogeneity in local data distributions often results in uneven model performance across clients Li et al. (2020); Wang et al. (2020). Since this outcome may be undesirable, a reasonable objective in such settings is to add constraints to ensure that the model’s loss is comparable across all clients.
We formulate the above problem as a constrained optimization problem as follows: consider solving the logistic regression for $C$ clients. For each client $c\in\{1,2,\dots,C\}$, it is associated with its own dataset $D_{c}=\{(x_{i},y_{i})\}_{i=1}^{|D_{c}|}$, where the label is $y_{i}\in\{-1,1\}$ and data feature is $x_{i}\in R^{d}$. For each client $c$, its own logistic regression loss $R_{c}(\theta)$ is defined as:

### v3

@@ 3.2 Relationship with SQP
$\dot{x}=-T(x)\Big(\nabla f(x)-J_{h}(x)^{\top}\left({J_{h}(x)T(x)J_{h}(x)^{\top}}\right)^{-1}(J_{h}(x)T(x)\nabla f(x)-Kh(x))\Big).$
@@
$x_{t+1}\!=\!x_{t}\!-\!\eta T(x_{t})\left(\!\nabla f(x_{t})\!-\!J_{h}(x_{t})^{\!\top}\!\left(J_{h}(x_{t})T(x_{t})J_{h}(x_{t})^{\!\top}\right)^{-1}\!(J_{h}(x_{t})T(x_{t})\nabla f(x_{t})\!-\!Kh(x_{t}))\right)$ (8)
@@
$x_{t+1}$ $=\textstyle\operatorname*{arg\,min}_{x}\nabla f(x_{t})^{\top}(x-x_{t})+\frac{1}{2\eta}(x-x_{t})^{\top}T(x_{t})^{-1}(x-x_{t})$
$\qquad\textstyle s.t.\quad h(x_{t})+J_{h}(x_{t})(x-x_{t})=0$ (9)
@@
Under Assumption 3, when $K=\frac{1}{\eta}I$, the discretization of FL (8) is equivalent to the SQP algorithm (3.2).
The proof of Theorem 2 leverages the fact that (3.2) satisfies the relaxed Slater condition. The detailed proof is deferred to Appendix D.
@@
Theorem 2 provides insights into the selection of $T(x)$ for the FL approach. Different choices of $T(x)$ lead to different types of SQP algorithms. Here we discuss two specific types of $T(x)$. First, $T(x)$ is set as the inverse of the Hessian matrix, i.e., $T(x)=\left(\nabla^{2}f(x)\right)^{-1}$, then (3.2) corresponds to the Newton-type algorithm where the quadratic term in the objective function is given by $(x-x_{t})^{\top}\nabla^{2}f(x)(x-x_{t})$, which is widely considered in literature (cf. Nocedal and Wright (2006); Bonnans et al. (2006)). For this specific type of $T(x)$, we name its corresponding FL dynamics (6) as the FL-Newton method. However, in the setting where the Hessian information is not available, another choice of $T(x)$ is simply setting it as the identity matrix $T(x)=I$, which is considered in recent works such as Oztoprak et al. (2023). In this case, the objective function resembles a proximal operator (cf. Boyd (2004); Parikh et al. (2014)), hence we name this as FL-proximal method. Due to space limit, we defer a more comprehensive overview of SQP to Appendix A. We would also like to emphasize that FL-proximal belongs to the class of first-order methods as its update only requires the first-order information $\nabla f(x),J_{h}(x)$.
Remark 4 (Comparison with other first-order methods).
To this end, we briefly compare FL-proximal methods with other first-order approaches, including Primal-Dual Gradient Descent (PDGD), Projected Gradient Descent (PGD), and the Augmented Lagrangian Method (ALM), with further details in Appendix A. PDGD has been well studied (Kose, 1956; Wang and Elia, 2011; Qu and Li, 2018), but is largely limited to convex settings and may fail in nonconvex problems (Zhang and Luo, 2020; Applegate et al., 2024). PGD also struggles with nonconvexity, as projections onto nonconvex sets are often intractable. ALM can handle nonconvex constraints, but each iteration requires solving a potentially expensive nonconvex subproblem. In contrast, when the constraint dimension is small relative to that of $x$, SQP-based methods like ours often perform better in practice Gould et al. (2003); Liu et al. (2018).
@@ 4 Extension to inequality constraints
$\begin{split}\textstyle-\nabla f(x)-J_{h}(x)^{\top}\lambda=0,\quad h(x)\leq 0,\quad\lambda\geq 0,\quad\lambda^{\top}h(x)=0\end{split}$ (11)
To measure convergence, we define the KKT-gap of the state variable $x$ and nonnegative control variable $\lambda\geq 0$ as follows:
$\textstyle\textup{{KKT-gap}}(x,\lambda):=\max\left\{\|\nabla f(x)+J_{h}(x)^{\top}\lambda\|,\left|\lambda^{\top}h(x)\right|,\max_{i}[h_{i}(x)]_{+}\right\},$
where $h_{i}(x)]_{+}=\max\{h_{i}(x),0\}$.
We can still view the problem as a control problem whose corresponding dynamics can be written as
$\begin{split}\textstyle\dot{x}=-T(x)\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right),\quad y=h(x),\quad\lambda\geq 0.\end{split}$ (12)
However, the problem becomes more complicated because we require the non-negativity constraints $\lambda\geq 0$ and complementary slackness $\lambda^{\top}h(x)=0$. At first glance, it is unclear how to guarantee these conditions through the control process. However, inspired by the relationship with SQP algorithms, we carefully design a more intricate FL controller as follows:
[…]
@@
Similar to the proof of Theorem 2, the proof of Theorem 3 also leverages strong duality and KKT conditions. The detailed proof is deferred to Appendix D.
@@
We define the index set $\mathcal{I}(x):=\{i:h_{i}(x)>0\}$. Our results rely on the following assumptions:
@@
[…]
Although Assumption 4 is quite complicated, there are some simplified versions that serve as a sufficient condition of Assumption 4. For example, if we start with a feasible $x(0)$, then $\mathcal{E}=\emptyset$ and hence Assumption 4 is automatically satisfied. Additionally, note that Assumption 3 is another sufficient condition of Assumption 4 (see Lemma 3 in Appendix G). Notably Assumption 4 is similar to the Mangasarian-Fromovitz constraint qualification (MFCQ) considered in literature Muehlebach and Jordan (2022a); Allibhoy and Cortés (2024)
@@
Let Assumption 1, 2 and 4 hold and let the control gain matrix $K$ be a diagonal matrix, i.e., $K=\textup{diag}\{k_{i}\}_{i=1},$ where $k_{i}>0$. Then the learning dynamics (13) satisfies the following properties
@@
    $\frac{dh_{i}(x(t))}{dt}\leq-k_{i}h_{i}(x(t))$, and hence the dynamic will converge to the feasible set.
@@
    $\textstyle\int_{t=0}^{T}\left(\|\nabla f(x(t))+J_{h}(x(t))\lambda(t)\|_{T(x(t))}^{2}-\sum_{i\in\mathcal{I}(x)^{c}}k_{i}\lambda_{i}(t)h_{i}(x(t))\right)dt\leq\ell(x(0))-\ell(x(T))$
@@
    $\inf_{0\leq t\leq T}\textup{{KKT-gap}}(x(t),\lambda(t))$ $\textstyle\leq\max\left\{\sqrt{\frac{2}{{\lambda_{\min}}T}\left(f(x(0))-f_{\min}+L\sum_{i\in\mathcal{I}(x(0))}h_{i}(x(0))\right)},\right$.
    $\textstyle\left.\frac{1}{\min_{i}k_{i}}\frac{2}{T}\left(f(x(0))-f_{\min}+(L+1)\sum_{i\in\mathcal{I}(x(0))}h_{i}(x(0))\right)\right\}\sim O\left(\frac{1}{\sqrt{T}}\right)$
    Further we have that KKT-gap asymptotically converges to zero, i.e. $\lim_{t\to+\infty}\textup{{KKT-gap}}(x(t),\lambda(t))=0$
Statement 4 in Theorem 4 implies that the algorithm can find an $\epsilon$-first-order KKT-point within time $\frac{1}{\epsilon^{2}}$. Similar to Theorem 1, the key step of the proof is to construct the merit function in Statement 2 (detailed proof deferred to Appendix E).
@@ 5 Momentum Acceleration for Constrained Optimization
$w_{t}=x_{t}+\beta(x_{t}-x_{t-1}),\quad x_{t+1}=w_{t}-\eta\nabla f(w_{t})$ (16)
Its corresponding continuous-time analogue can be written as a second-order ODE Polyak (1964); Su et al. (2016)
$\dot{x}=z,\quad\dot{z}=-\alpha z-\nabla f(x)$ (17)
@@
$\textstyle w_{t}=x_{t}+\beta(x_{t}-x_{t-1}),\quad x_{t+1}=w_{t}+\eta\nabla f(w_{t})+J_{h}(w_{t})^{\top}\lambda_{t}$
$\textstyle\lambda_{t}=-\left(J_{h}(w_{t})J_{h}(w_{t})^{\top}\right)^{-1}(J_{h}(w_{t})\nabla f(w_{t})-\frac{1}{\eta}h(w_{t}))$ (18)
@@
FL-momentum for Equality-Constrained Optimization $\begin{split}\dot{x}&=z\qquad\dot{z}=-\alpha z-\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)\\ \lambda&=-(J_{h}(x)J_{h}(x)^{\top})^{-1}(J_{h}(x)\nabla f(x)-Kh(x))\end{split}$ (19)
Note that compared with FL-proximal (8), the difference in (5) is the addition of a momentum step $w_{t}=x_{t}+\beta(x_{t}-x_{t-1})$. Similarly we can propose the FL-momentum scheme for inequality constraint case as follows:
FL-momentum for Inquality-Constrained Optimization $\dot{x}$ $=z,\qquad\dot{z}=-\alpha z-\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)$ (20) $\lambda$ $=\textstyle\operatorname*{arg\,min}_{\lambda\geq 0}\frac{1}{2}\lambda^{\top}J_{h}(x)J_{h}(x)^{\top}\lambda+\lambda^{\top}\left(J_{h}(x)\nabla f(x)-Kh(x)\right)$
The numerical simulation in Section 6 (Figure 2) suggests that momentum methods indeed accelerate the convergence rate. We would also like to note that as far as we know, the acceleration of SQP methods are generally achieved via Newton or quasi-Newton methods, there’s little work on exploring acceleration via momentum approaches, which makes our proposed momentum algorithm a novel contribution.
5.1 Convergence Analysis for FL-momentum: Nonconvex Case
@@
Apart from Assumption 1 and 2, we also make the following assumptions on $f$ and $h$.
@@
Both $f(x)$, $h(x)$ are three-times differentiable and the derivatives are bounded, thus, we know that there exist some constants $L_{f},L_{1},L_{2}$ such that
$\textstyle\|\nabla^{2}f(x)\|\leq L_{f},~~\left\|\frac{\partial\bar{\lambda}(x)}{\partial x}\right\|\leq L_{1},~~\left\|\frac{\partial\left(J_{h}(x)^{\top}\lambda(x)+\left(\frac{\partial\bar{\lambda}(x)}{\partial x}^{\top}h(x)\right)\right)}{\partial x}\right\|\leq L_{2}$
@@
We also assume that that there exists a constant $\bar{H}$ such that $\|\bar{H}(x)\|\leq\bar{H},~~\forall x,$ where $\bar{H}(x):=[h(x)^{\top}\nabla^{2}h_{i}(x)]_{i=1}^{n}$.
@@
Assume that Assumption 1, 2, 5 and 6 hold. Let two positive constants $a_{1},a_{2}$ be such that $a_{2}\geq\left(4\frac{\lambda_{\max}(K)}{{\lambda_{\min}}(K)}L_{2}D+\frac{L_{1}^{2}}{{\lambda_{\min}}(K)}\right)\times a_{1}\geq 0.$ We define the following merit function:
@@
then for $\alpha\geq\left(a_{1}(L_{f}+L_{2})+a_{2}(M^{2}+\bar{H})+\frac{1}{a_{1}}+\frac{2(\lambda_{\max}(K)D^{2})}{a_{2}}\right)+1$, we have that
@@
    $\hskip-20.0pt\int_{t=0}^{T}\frac{a_{2}{\lambda_{\min}}(K)}{8}\|h(x(t))\|^{2}+\frac{a_{1}}{4}\|\nabla f(x(t))+J_{h}(x(t))^{\top}\overline{\lambda}(x(t))\|^{2}dt\leq\ell(x(0),z(0))-\min_{x,z}\ell(x,z)$
@@
    $\textstyle\quad\inf_{0\leq t\leq T}\textup{{KKT-gap}}(x(t),\overline{\lambda}(x(t)))\leq\sqrt{\frac{\ell(x(0),z(0))-\ell_{\min}}{\min\left\{\frac{a_{2}{\lambda_{\min}}(K)}{8},\frac{a_{1}}{4}\right\}T}}\sim O(\frac{1}{\sqrt{T}})$
    and $\lim_{t\to+\infty}\textup{{KKT-gap}}(x(t),\overline{\lambda}(x(t)))=0,\quad\textstyle\lim_{t\to+\infty}\textup{{KKT-gap}}(x(t),\lambda(x(t)))=0.$
The detailed proof is provided in Appendix F.
Remark 5 (Limitation of the result).
One limitation of Theorem 5 is that it establishes convergence but not acceleration over FL-proximal. However, when the constraint function $h(x)$ is affine, the algorithm is equivalent to the momentum-accelerated projected gradient method (see Appendix F.1), offering insight into its potential for accelerating optimization.
6 Numerical Verifications
For numerical validation, we consider a logistic regression problem involving heterogeneous clients Shen et al. (2022); Hounie et al. (2024). Many scenarios, such as federated learning and fair machine learning, require training a common model in a distributed manner by utilizing data samples from diverse clients or distributions. In practice, heterogeneity in local data distributions often results in uneven model performance across clients Li et al. (2020); Wang et al. (2020). Since this outcome may be undesirable, a reasonable objective in such settings is to add constraints to ensure that the model’s loss is comparable across all clients.
We formulate the above problem as a constrained optimization problem as follows: consider solving the logistic regression for $C$ clients. For each client $c\in\{1,2,\dots,C\}$, it is associated with its own dataset $D_{c}=\{(x_{i},y_{i})\}_{i=1}^{|D_{c}|}$, where the label is $y_{i}\in\{-1,1\}$ and data feature is $x_{i}\in R^{d}$. For each client $c$, its own logistic regression loss $R_{c}(\theta)$ is defined as